# AI-assisted Programming May Decrease the Productivity of Experienced Developers by Increasing Maintenance Burden

### v2

@@
AI-assisted Programming May Decrease the Productivity of Experienced Developers by Increasing Maintenance Burden
@@
Generative AI solutions like GitHub Copilot have been shown to increase the productivity of software developers. Yet prior work remains unclear on the quality of code produced and the challenges of maintaining it in software projects. If quality declines as volume grows, experienced developers face increased workloads reviewing and reworking code from less-experienced contributors. We analyze developer activity in Open Source Software (OSS) projects following the introduction of GitHub Copilot. We find that productivity indeed increases. However, the increase in productivity is primarily driven by less-experienced (peripheral) developers. We also find that code written after the adoption of AI requires more rework. Importantly, the added rework burden falls on the more experienced (core) developers, who review 6.5% more code after Copilot’s introduction, but show a 19% drop in their original code productivity. More broadly, this finding raises caution that productivity gains of AI may mask the growing burden of maintenance on a shrinking pool of experts.
@@
Generative AI, GitHub Copilot, Open Source Software, Software Maintenance, Software Quality, Difference-in-Differences
@@
How will AI shape the future of software development? This question has taken on renewed significance with the recent rise of generative AI (GenAI) technologies, which are becoming an integral part of daily operations of software development teams. A prominent example is GitHub Copilot, an AI-powered coding assistant designed to support developers by generating code suggestions and accelerating routine programming tasks. When GitHub launched Copilot, it was introduced as “your AI pair programmer," emphasizing not only its role as an automation tool but also as a team member who partners with the developer to write and test code (Friedman 2021). Unlike earlier code automation tools, the framing of GitHub Copilot as a pair programmer suggests that the future of software development will increasingly involve Human-AI pair programming.
@@
For organizations and communities involved in software development, the addition of AI pair programmers in teams offers the potential for significant productivity gains. Recent research shows that developers who use GitHub Copilot complete their programming tasks 55.8% faster (Peng et al. 2023). Such productivity benefits lead to promises of faster time-to-market and increased revenue for organizations developing software applications. Considering this, major tech organizations have started to increasingly rely on AI in their projects - “more than a quarter of all new code at Google is generated by AI, then reviewed and accepted by engineers," reported Google CEO Sundar Pichai in January, 2025.111https://www.technologyreview.com/2025/01/20/1110180/the-second-wave-of-ai-coding-is-here/ While these productivity gains are promising, they also raise important questions about the quality and maintainability of AI-generated code. Because AI tools can lower the skill barrier for writing code (Dakhel et al. 2023), developers may increasingly rely on them without fully understanding the implications of the code being produced (Barrett et al. 2023). This growing reliance introduces new risks such as security vulnerabilities, bugs, and suboptimal solutions that may undermine the long-term stability and maintainability of software systems.
The potential challenges that AI poses to software quality and security are expected to be especially pronounced in distributed software development teams, such as in Open Source Software (OSS) communities. In these communities, contributors from around the world collaborate, often voluntarily, to develop and maintain software that form the digital infrastructure of our society (e.g., Linux, Apache, LaTeX, Python), making it freely or cheaply available to the public (Eghbal 2020, Nagle 2019). Despite the voluntary nature of work in these communities, the value of OSS (the estimated cost for firms to build equivalent software internally) is estimated at \$8.8 trillion (Hoffmann et al. 2024).  [^2] These estimates suggests that firms would spend approximately 3.5 times more on software than they currently do if OSS did not exist (Hoffmann et al. 2024). Importantly, OSS is not only a source of low cost software but also a model that many firms now adopt in their own development practices. Microsoft, once a vocal critic of open source, has become a key advocate—open-sourcing its core technologies such as the .NET Framework, TypeScript, PowerShell, and VSCode since the mid-2010s. [^3] https://opensource.microsoft.com/ This shift reflects a broader industry trend toward embracing OSS as a development model for their software applications (Nagle 2019).
Given this critical role for both firms and society, the growing adoption of AI in OSS communities raises important questions about its broader impact. On the one hand, AI tools can lower the barrier for peripheral contributors (contributors who come from the community of users of software; Rullani and Haefliger 2013) to contribute to these software projects. On the other hand, this surge in AI-assisted contributions may introduce new maintenance burdens, particularly around code quality and security (Eghbal 2020). In this study, we examine how AI adoption is reshaping the OSS development workflow by introducing new maintenance challenges, ultimately raising concerns about the long-term sustainability and reliability of the digital infrastructure that our society increasingly depends on.
The quality maintenance of OSS communities is typically carried out by a small group of experienced (core) contributors who take on the role of gatekeepers of the community. They play a crucial role in ensuring the project’s quality, stability, and sustenance by reviewing contributions, managing releases, and guiding the community (Eghbal 2020). These contributors are deeply embedded in the community and often take on the “less desirable" maintenance related activities that are necessary for the upkeep of the community (Eghbal 2020, Nagle et al. 2020, Nagle 2019). While these core contributors are critical for the OSS community, they are often few, under-incentivized and overworked (Eghbal 2020). Consequently, the current state of much of the OSS infrastructure is characterized by an over-reliance on the efforts of a few core contributors to perform all the maintenance work (Osborne 2024). For example, the backbone of many data science applications is the Python module “pandas”, which is used by over 2.7 million users. The “pandas” OSS project has been developed with contributions from more than 3,400 contributors who have helped build its features over time. [^4] https://github.com/pandas-dev/pandas Despite its wide adoption, the project maintenance has rested on the shoulders of only about 35 core contributors, of whom just 15 are currently active. [^5] https://pandas.pydata.org/about/team.html [^6] This point was raised in a keynote by Fernando Perez, the founder of IPython/Jupyter, at the Euro SciPy conference in 2011, where he remarked that - https://numfocus.org/blog/why-is-numpy-only-now-getting-funded “Python in science is a great success story, but the entire edifice rests on (often the spare time of)  30 people.” Now, considering this, imagine a surge of new contributors submitting code to these projects (which might come from the productivity increase brought by AI), requiring the attention of the already overburdened maintainers. An increased workload could lead to delays in reviewing and merging the submitted code, greater difficulty in validating code quality, and potentially a higher risk of introducing vulnerabilities into the critical digital infrastructure.
To understand the impact of the adoption of AI pair programming on the maintenance of OSS communities, we examined how the OSS development workflow changes before and after the deployment of GitHub Copilot. The development workflow of an OSS project is illustrated in Figure 1. A feature of this development workflow is that it allows multiple contributors to work independently on separate branches (or forks) and submit their changes for inclusion in the main project branch. This is done through a pull request (PR), which serves as a formal proposal to merge code. Core or experienced contributors often handle key maintenance tasks: they review submitted PRs for quality and compliance, suggest improvements or corrections, and integrate approved changes into the main project. This workflow enables distributed development and community-driven innovation, allowing peripheral contributors who come from the community of users of the project to contribute. At the same time, it ensures that the code is reviewed, refined, and improved before merging into the main branch. The effectiveness of this review process has helped OSS achieve remarkably high quality, surpassing proprietary software in metrics like bugs per 1,000 lines of code. [^7] https://blogs.worldbank.org/en/opendata/quality-open-source-software-how-many-eyes-are-enough However, the success and effectiveness of this "decentralized-development" approach hinges on the core contributors who actively participate in the review-rework process of the PRs submitted by different contributors (Rullani and Haefliger 2013).
Figure 1: The workflow of OSS projects. It comprises of a primary branch of a GitHub (project), which typically contains the main source code that serves as the foundation for new feature development, bug fixes, and updates. Changes to this branch are usually controlled through a structured review process conducted by maintainers to ensure code quality and prevent issues. To contribute to the project, a PR is submitted to propose changes to a project. It allows developers to submit modifications, request feedback, and merge updates into the main branch. The review and refinement process comprises of two activities that we measure in our study - pull request review (PR Review) and pull request rework (PR Rework). Eventually the reviewed and reworked changes will either be merged into the main branch or undergo additional rounds of review and rework.
@@
Our main objective is to examine whether the review and rework effort on PRs changed after the introduction of GitHub Copilot. To empirically test this, we exploit the release of GitHub Copilot as a technical preview in June 2021, which included limited programming language endorsement. We focus on OSS projects owned by Microsoft, as the company had exclusive access to OpenAI’s GPT-3, the model powering GitHub Copilot during its technical preview, due to its investment in OpenAI and its prior acquisition of GitHub. [^8] https://www.technologyreview.com/2020/09/23/1008729/openai-is-giving-microsoft-exclusive-access-to-its-gpt-3-language-model/ [^9] https://www.mckinsey.com/industries/technology-media-and-telecommunications/our-insights/thomas-dohmke-on-improving-engineering-experience-using-generative-ai The individual users in our dataset are contributors to Microsoft-owned OSS projects. We estimate the effect of Copilot at both the project and contributor levels using a Difference-in-Differences (DiD) design. Treatment and control groups were defined based on the primary programming language: those using Copilot-endorsed languages formed the treatment group, while non-endorsed language users served as the control (Yeverechyahu et al. 2024). For both project and contributor levels, we collected data on programming activities and aggregated them at the monthly level.

### v3

# AI-Assisted Programming Decreases the Productivity of Experienced Developers by Increasing the Technical Debt and Maintenance Burden

@@
AI-Assisted Programming Decreases the Productivity of Experienced Developers by Increasing the Technical Debt and Maintenance Burden
@@
GenAI solutions like GitHub Copilot have been shown to increase the productivity of software developers. Yet prior work remains unclear on the quality of code produced and the challenges of maintaining it in software projects. If quality declines as volume grows, technical debt accumulates as experienced developers face increased workloads reviewing and reworking code from less-experienced contributors. We analyze developer activity in Open Source Software (OSS) projects following the introduction of GitHub Copilot. We find that productivity indeed increases. However, the increase in productivity is primarily driven by less-experienced (peripheral) developers. We also find that code written after the adoption of AI requires more rework to satisfy repository standards, indicating a potential increase in technical debt. Importantly, the added rework burden falls on the more experienced (core) developers, who review 6.5% more code after Copilot’s introduction, but show a 19% drop in their original code productivity. More broadly, this finding raises caution that productivity gains of AI may mask the growing burden of maintenance on a shrinking pool of experts, together with increased technical debt for the projects. The results highlight a fundamental tension in AI-assisted software development between short-term productivity gains and long-term system sustainability.
@@
GenAI, GitHub Copilot, Open Source Software, Software Maintenance, Technical Debt, Difference-in-Differences
1 Introduction
@@
How will AI shape the future of knowledge-intensive industries? This question has taken on renewed significance with the recent rise of Genarative AI (GenAI) technologies, which are becoming an integral part of daily operations of software development, scientific research, healthcare and other expert-driven fields. A prominent example is GitHub Copilot, an AI-powered coding assistant designed to support developers by generating code suggestions and accelerating routine programming tasks (Peng et al. 2023). When GitHub launched Copilot, it was introduced as “your AI pair programmer," emphasizing not only its role as an automation tool but also as a team member who partners with the developer to create knowledge (Friedman 2021). Unlike earlier coding automation tools that primarily targeted productivity, GitHub Copilot’s framing as a pair programmer signals a deeper shift. It implies that AI may fundamentally reshape how knowledge-intensive work is performed, coordinated, and organized, rather than merely accelerating existing tasks.
@@
For organizations and communities involved in software development, the addition of AI pair programmers in teams offers the potential for significant productivity gains. Right after the launch of GitHub Copilot, research shows that developers who use Copilot completed their programming tasks 55.8% faster (Peng et al. 2023). Such productivity benefits lead to promises of faster time-to-market and increased revenue for organizations developing software applications. Considering these shifts, major tech organizations have started to increasingly rely on AI in their projects - “more than a quarter of all new code at Google is generated by AI, then reviewed and accepted by engineers," reported Google CEO Sundar Pichai in January, 2025.111https://www.technologyreview.com/2025/01/20/1110180/the-second-wave-of-ai-coding-is-here/ Moreover, Microsoft CTO Kevin Scott expects that 95% of all code will be AI-generated by 2030.222https://www.businessinsider.com/microsoft-cto-ai-generated-code-software-developer-job-change-2025-4
While these productivity gains are promising, they also raise important questions about the quality and maintainability of AI-generated code. Because AI tools can lower the skill barrier for writing code (Dakhel et al. 2023), AI tools enable broader participation but may also encourage developers to rely on generated solutions without fully understanding the underlying design rationale and potential integration issues (Barrett et al. 2023). Such reliance increases the likelihood of quick fixes that favor short-term functionality over long-term maintainability (Barrett et al. 2023). Extant literature characterizes “quick and dirty” software customizations made without a complete understanding of their future implications as technical debt, as they undermine system reliability and impose long-term maintenance obligations (Kruchten et al. 2012, Brown et al. 2010, Banker et al. 2021). As a result, project maintainers must devote additional effort to understanding, reviewing, and reworking AI-generated code before it can be safely integrated. In our context, we contend that the growing reliance on AI-assisted development may accelerate the accumulation of technical debt, as design shortcuts taken to expedite system deployment become embedded in software systems (Ramasubbu and Kemerer 2016, 2021).
The technical debt and maintenance challenges that AI poses are expected to be especially pronounced in distributed software development teams, such as in Open Source Software (OSS) communities. In these communities, contributors from around the world collaborate, often voluntarily, to develop and maintain software that form the digital infrastructure of our society (e.g., Linux, Apache, LaTeX, Python), making it freely or cheaply available to the public (Eghbal 2020, Nagle 2019). Despite the voluntary nature of work in these communities, OSS constitutes critical digital infrastructure for modern society, with estimates suggesting that the total cost of reproducing this software would amount to \$8.8 trillion (Hoffmann et al. 2024). [^3] These estimates suggests that firms would spend approximately 3.5 times more on software than they currently do if OSS did not exist (Hoffmann et al. 2024). Given this critical role for both firms and society, the growing adoption of AI in OSS communities raises important questions about its broader impact. On the one hand, AI tools can lower the barrier for peripheral contributors (relatively newer contributors to the project who come from the community of users of software; Rullani and Haefliger 2013) to contribute to these software projects. On the other hand, this surge in AI-assisted contributions may result in a build up of technical debt and increase the maintenance obligations of the software (Banker et al. 2021). If this mechanism holds, the performance consequences of technical debt are likely to intensify over time, as mounting maintenance burdens increasingly divert expert attention away from productive development activities, thereby amplifying the negative impact of technical debt on project and firm performance.
This study seeks to address this tension. We can distinguish two important types of activities in OSS projects - development and maintenance. The development activity in OSS projects involves producing code and submitting pull requests (PRs) to the project. The open nature of OSS communities allows anyone—regardless of skill or experience - to contribute to this process (Rullani and Haefliger 2013). In contrast, maintenance involves reactive tasks such as triaging issues, reviewing submitted PRs, suggesting modifications or corrections, and ensuring that contributions align with project standards (Banker et al. 2021, Ramasubbu and Kemerer 2021). These maintenance tasks are generally less intrinsically driven and typically assigned to contributors who possess both technical expertise and trust of the community to ensure the quality and reliability of the codebase (Medappa et al. 2023, Eghbal 2020). Thus, we seek to answer the question: How does AI-assisted programming impact the development and maintenance activities and the technical debt of the OSS projects?
@@
In this study, we examine whether technical debt and maintenance efforts of OSS projects changed after the introduction of GitHub Copilot through increased code review and rework effort on PRs. To empirically test this, we exploit the release of GitHub Copilot as a technical preview in June 2021, which included limited programming language endorsement. We focus on OSS projects owned by Microsoft, as the company had exclusive access to OpenAI’s GPT-3, the model powering GitHub Copilot during its technical preview, due to its investment in OpenAI and its prior acquisition of GitHub. [^4] https://www.technologyreview.com/2020/09/23/1008729/openai-is-giving-microsoft-exclusive-access-to-its-gpt-3-language-model/ [^5] https://www.mckinsey.com/industries/technology-media-and-telecommunications/our-insights/thomas-dohmke-on-improving-engineering-experience-using-generative-ai The individual users in our dataset are contributors to Microsoft-owned OSS projects. We estimate the effect of Copilot at both the project and contributor levels using a Difference-in-Differences (DiD) design. Treatment and control groups were defined based on the primary programming language: those using Copilot-endorsed languages formed the treatment group, while non-endorsed language users served as the control (Yeverechyahu et al. 2024). For both project and contributor levels, we collected data on programming activities and aggregated them at the monthly level.
We examine the changes in code productivity after Copilot by three measures: lines of code added, commits  [^6] A commit is the fundamental unit of change on GitHub. Similar to saving a file that’s been edited, a commit records changes to one or more files on GitHub - https://docs.github.com/en/pull-requests/committing-changes-to-your-project/creating-and-editing-commits/about-commits and PRs submitted to the project. To capture the secondary effects of AI adoption on technical debt and maintenance effortsof OSS communities, we focused on two complementary outcomes. First, we measured technical debt at the project level using PR rework, which reflects the extent to which initially submitted contributions require modification before being integrated into the codebase (Ramasubbu and Kemerer 2016, 2021). Higher levels of rework indicate greater reliance on expedient or insufficiently integrated solutions. Second, we measured maintenance effort at the individual level using PR reviews, which capture the time and effort required from contributors to evaluate, correct, and integrate submitted code (Medappa et al. 2023).